OpenStudy (anonymous):

The sum of two consecutive odd integers is 236. What is the smaller integer?

OpenStudy (anonymous):

4x = 236 x = 59 --- 1st: 2x-1 = 117 2nd: 2x+1 = 119

OpenStudy (anonymous):

tadaa! any more questions?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

In a certain county, the number of Charter Schools is eight less than twice the number of Alternative Schools. We know that there are 36 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (anonymous):

Let the odd integer be : x So, the other will be: (x + 2) x + x +2 = 236 2x = 234 Find x from here and get the answer...

OpenStudy (anonymous):

lemme show u how igot it

OpenStudy (anonymous):

Let the Alternative Schools be x Number of Charter School = 2x - 8 So, 2x - 8 = 36 So, x = 44 x = 22 is the answer...

OpenStudy (anonymous):

tyyyyyy

OpenStudy (anonymous):

what about this

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter?

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

Let his Daughter's age be x 69 = 4x - 3 4x = 72 x = 18

OpenStudy (anonymous):

lol ty one more!

OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

The answer goes like this: Let Sam's age is x So, Coach age is : 2x Before 10 years ago, Sam is (x - 10) And Coach was (2x - 10) So, 2x - 10 = 4(x - 10) x - 5 = 2x - 20 x = 15...

OpenStudy (anonymous):

4x - 4 = 2x + 4 + 6 2x = 14 and get x from here...

OpenStudy (anonymous):

uh...

OpenStudy (anonymous):

What happened??

OpenStudy (anonymous):

i dident get it

OpenStudy (anonymous):

Distribute 4 and 2 to the brackets.. What you get???

OpenStudy (anonymous):

Solve them: 4(x - 1) = ???? And, 2(x + 2) = ???? Tell me what will you get????

OpenStudy (anonymous):

Ugh... idk

OpenStudy (anonymous):

Then, you must know that...

OpenStudy (anonymous):

:(